# S3E: A Mulit-Robot Multimodal Dataset for Collaborative SLAM

### v7

@@ I Introduction
Collaborative Simultaneous Localization and Mapping (C-SLAM) is crucial for multi-agent cooperation, improving the robustness and efficiency of localization and mapping in shared spaces [1, 2, 3, 4]. Despite progress, repeatability and benchmarking in C-SLAM research face challenges due to: 1) System Complexity: C-SLAM systems’ intricate software and sensor integration make perfect replication difficult, highlighting the need for public datasets for methodological evaluation. 2) Evaluation Methodology: Traditional approaches, which divide single-agent SLAM datasets into segments for virtual agents, are impractical for real-world scenarios due to uniform perspectives and lighting at intersection points [5]. These limitations underscore the importance of developing more effective evaluation strategies and datasets to advance C-SLAM research.
Furthermore, effective trajectory design for multi-robot scenarios in C-SLAM must follow two key principles: 1) Temporal and Spatial Diversity: Trajectories should be crafted to provide varying intra- and inter-loop closures at different times and places, ensuring robots have complementary perspectives. 2) Communication Constraints: Robots are typically limited to sharing information within close proximity, necessitating trajectory designs that maintain a reasonable interaction distance for effective communication. In this paper, we introduce four trajectory prototypes designed to meet these principles and evaluate the adaptability of C-SLAM methodologies across diverse closure strategies in multi-robot operations.
@@ II Related Work
Multi-Agent SLAM Datasets. The UTIAS dataset [8] was a pioneering work with five robots exploring a small area. The AirMuseum dataset [10] introduced a heterogeneous multi-robot setup in a warehouse, and the FordAV dataset [9] presented a 3D map with multiple autonomous vehicles capturing seasonal changes. However, these datasets often have significant co-view overlap, limiting comprehensive C-SLAM evaluation.
@@ III-A Sensor Configuration
The technical specifications of the sensors and ground truth devices integrated into our payload are outlined on our project website [^1] https://pengyu-team.github.io/S3E. This includes the sensor types, their resolution, measurement range, accuracy, and any other pertinent technical details that define their contribution to the SLAM system’s performance.
@@ III-D Trajectory Paradigms
Env. Region a b c d Alpha Bob Carol Camera IMU LiDAR UWB
@@ III-F Dataset Format
Our research utilizes the ROS2 [14] bag format for sensor data storage, a standard in robotics known for efficient data management and playback. To simplify data replay and ensure easy access to synchronized sensor streams from our Alpha, Blob, and Carol robots, we’ve implemented a strategy to merge data from the same operational sequence into one ROS2 bag file.
IV Experiments
Table V: ATE $[m]$ for Single SLAM and C-SLAM in the S3Ev1.0 outdoor environment without UWB measurement. $\alpha$, $\beta$, and $\gamma$ stand for Alpha, Bob, and Carol respectively. ✗ fails to initialize or track frames. If inter-loop closures detection fails, we mark it “Failed”.
@@ IV Experiments
Table VI: ATE $[m]$ for Single SLAM and C-SLAM in the S3Ev1.0 indoor environment without UWB measurement. $\alpha$, $\beta$, and $\gamma$ stand for Alpha, Bob, and Carol respectively.
@@
Table VII: ATE $[m]$ for C-SLAM in the S3Ev2.0 outdoor environment with UWB measurement. $\alpha$, $\beta$, and $\gamma$ stand for Alpha, Bob, and Carol respectively.
@@ IV-B Results
V Conclusion

### v8

# S3E: A Multi-Robot Multimodal Dataset for Collaborative SLAM

@@ Abstract
Index Terms:
Multi-Robot SLAM; Data Sets for SLAM; SLAM
@@ I Introduction
Collaborative Simultaneous Localization and Mapping (C-SLAM) is crucial for multi-agent cooperation, improving the robustness and efficiency of localization and mapping in shared spaces [1, 2, 3, 4]. Despite progress, repeatability and benchmarking in C-SLAM research face challenges due to: 1) System Complexity: The complexity of C-SLAM systems, which integrate intricate software architectures and a variety of hardware, complicates the process of achieving perfect replication, thereby emphasizing the need for publicly available datasets to facilitate methodological evaluation. 2) Evaluation Methodology: Traditional approaches, which divide single-agent SLAM datasets into segments for virtual agents, are impractical for real-world scenarios due to uniform perspectives and lighting at intersection points [5]. These limitations underscore the importance of developing more effective evaluation strategies and datasets to advance C-SLAM research.
Furthermore, effective trajectory design for multi-robot scenarios in C-SLAM must follow two key principles: 1) Temporal and Spatial Diversity: Trajectories should be crafted to provide multiple intra- and inter-loop closures at different times and places along the trajectories, ensuring robots have complementary perspectives. 2) Communication Constraints: Robots are typically limited to sharing information within close proximity, necessitating trajectory designs that maintain a reasonable interaction distance for effective communication. In this paper, we introduce four trajectory prototypes designed to meet these principles and evaluate the adaptability of C-SLAM methodologies across diverse closure strategies in multi-robot operations.
@@ II Related Work
Multi-Agent SLAM Datasets. The UTIAS dataset [8] was a pioneering work with five robots exploring a small area. The AirMuseum dataset [10] introduced a heterogeneous multi-robot setup in a warehouse, and the FordAV dataset [9] is collected by a fleet of autonomous vehicles, providing a multi-agent perspective with seasonal variations in weather, lighting, construction, and traffic conditions. However, these datasets often have significant co-view overlap, limiting comprehensive C-SLAM evaluation.
@@ III-A Sensor Configuration
The technical specifications of the sensors and ground truth devices integrated into our payload are outlined on our project website [^1] https://pengyu-team.github.io/S3E. This includes the sensor types, their resolution, measurement range, accuracy, and any other pertinent technical details that define their contribution to the SLAM system’s performance. In particular, we utilize the centermeter-accuracy GNSS and RTK for outdoor environments, achieving a precision of ±$\pm 1\text{cm}$, to record the ground truth trajectories. For indoor environments, where GNSS signals are typically unavailable, we utilize a millimeter-precision motion capture system to obtain our ground truth data.
@@ III-D Trajectory Paradigms
Env. Region a b c d Alpha ($\alpha$) Bob ($\beta$) Carol ($\gamma$) Camera IMU LiDAR UWB
@@ III-F Dataset Format
Table V: ATE $[m]$ for Single SLAM and C-SLAM in the S3Ev1.0 outdoor environment without UWB measurement. ✗ fails to initialize or track frames. If inter-loop closures detection fails, we mark it “Failed”.
@@
Table VI: ATE $[m]$ for Single SLAM and C-SLAM in the S3Ev1.0 indoor environment without UWB measurement.
@@
Table VII: ATE $[m]$ for C-SLAM in the S3Ev2.0 outdoor environment with UWB measurement.
@@
Our research utilizes the ROS2 [14] bag format for sensor data storage, a standard in robotics known for efficient data management and playback. To simplify data replay and ensure easy access to synchronized sensor streams from our Alpha ($\alpha$), Blob ($\beta$), and Carol ($\gamma$) robots, we’ve implemented a strategy to merge data from the same operational sequence into one ROS2 bag file.
IV Experiments
IV-A Baselines
@@ IV-B Results
V Known Issues
The S3E dataset, while a substantial contribution to the field of C-SLAM, presents certain limitations. This section delves into the known issues associated with the dataset and the broader challenges of C-SLAM systems:
  1. 1.
    Scalability: C-SLAM systems must efficiently handle operations with a large number of robots without significant performance degradation.
  2. 2.
    Resource Management: There are strict limitations on communication and computation capabilities of mobile robots, which C-SLAM must overcome to achieve real-time performance.
VI Conclusion